OpenStudy (anonymous):

Verify each identity: (sect+1)/tant=tant/(sect-1)

OpenStudy (anonymous):

\[\left( \frac{ \sec t+1 }{ \tan t } \right)*\frac{ \tan t }{ \tan t }=\frac{ \tan t \left( \sec t+1 \right) }{\tan ^{2}t }\] \[=\frac{ \tan t \left( \sec t+1 \right) }{\sec ^{2}t-1 }=\frac{ \tan t \left( \sec t+1 \right) } {\left( \sec t+1 \right)\left( \sec t-1 \right) }\] =?

OpenStudy (anonymous):

thank you!!!!!!!!!!!! :DDDDDDDDD

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@alessandro96